Potassium hydroxide is used to precipitate each of the cations from their respective solution. Determine the minimum concentration of KOH required for precipitation to begin in each case.

Part A

1.9×10−2 M CaCl2

Express your answer using two significant figures.

Part B

2.8×10−3 M Fe(NO3)2

Express your answer using two significant figures.

Part C

2.0×10−3 M MgBr2

Express your answer using two significant figures.

Thanks
Answer #1

A) The balanced chemical equation for the reaction between CaCl2 and NaOH is given as

CaCl2 (aq) + 2 KOH (aq) ---------> Ca(OH)2 (s) + 2 KCl (aq)

Ca(OH)2 (s) <=======> Ca2+ (aq) + 2 OH- (aq)

Ksp = [Ca2+][OH-]2 = 5.02*10-6

(http://www2.chm.ulaval.ca/gecha/chm1903/6_solubilite_solides/solubility_products.pdf)

Therefore,

5.02*10-6 = (1.9*10-2)*[OH-]2

(ignore units since Ksp is dimensionless)

=====> [OH-]2 = (5.02*10-6)/(1.9*10-2) = 2.642*10-4

=====> [OH-] = 0.0162 ≈ 0.016 (correct to 2 sig. figs)

The minimum concentration of KOH required to precipitate 1.9*10-2 M CaCl2 as Ca(OH)2 is 0.016 M (ans).

B) The balanced chemical equation for the reaction between CaCl2 and NaOH is given as

Fe(NO3)2 (aq) + 2 KOH (aq) ---------> Fe(OH)2 (s) + 2 KNO3 (aq)

Fe(OH)2 (s) <=======> Fe2+ (aq) + 2 OH- (aq)

Ksp = [Fe2+][OH-]2 = 4.87*10-17

(http://www2.chm.ulaval.ca/gecha/chm1903/6_solubilite_solides/solubility_products.pdf)

Therefore,

4.87*10-17 = (2.8*10-3)*[OH-]2

(ignore units since Ksp is dimensionless)

=====> [OH-]2 = (4.87*10-17)/(2.8*10-3) = 1.739*10-14

=====> [OH-] = 1.318*10-7 ≈ 1.3*10-7 (correct to 2 sig. figs)

The minimum concentration of KOH required to precipitate 2.8*10-3 M Fe(NO3)2 as Fe(OH)2 is 1.3*10-7 M (ans).

C) The balanced chemical equation for the reaction between CaCl2 and NaOH is given as

MgBr2 (aq) + 2 KOH (aq) ---------> Mg(OH)2 (s) + 2 KBr (aq)

Mg(OH)2 (s) <=======> Mg2+ (aq) + 2 OH- (aq)

Ksp = [Mg2+][OH-]2 = 5.61*10-12

(http://www2.chm.ulaval.ca/gecha/chm1903/6_solubilite_solides/solubility_products.pdf)

Therefore,

5.61*10-12 = (2.0*10-3)*[OH-]2

(ignore units since Ksp is dimensionless)

=====> [OH-]2 = (5.61*10-12)/(2.0*10-3) = 2.805*10-9

=====> [OH-] = 5.296*10-5 ≈ 5.3*10-5 (correct to 2 sig. figs)

The minimum concentration of KOH required to precipitate 2.0*10-3 M MgBr2 as Mg(OH)2 is 5.3*10-5 M (ans).